OpenStudy (anonymous):

y varies inversely as three times x. When x = 6, y = 2. Find x when y = 12. x = 1 x = 3 x = 6 x = 9

OpenStudy (anonymous):

Inverse variation can be written as y=k/x. k stands for "constant" (crazy, huh), so it doesn't change. Plug in what you're given and you get 2=k/(3*6) . Can you solve that for k?

OpenStudy (anonymous):

In case your wondering, you get k/(3*6) because y varies inversely as 3*x and x is given as 6. That's why you have 3*6 in the denominator.

OpenStudy (anonymous):

I know, you're waiting me out. OK, you win! So you have 2=k/18. Multiply both sides by 18 to eliminate the 18 on the right side and you get k=36. Now that you know what k (the constant) is, just plug it in to the new equation. Since the equation is y=k/3x and we know that y=12 and k=36, just plug those values into the equation for y and k and you'll have your answer! Let me know what you get.